OpenStudy (andertj21):

Which of the following equations is of a parabola with a vertex at (1, 1)?

OpenStudy (andertj21):

Its not a answer choice.... y = (x - 1)2 - 1 y = (x - 1)2 + 1 y = (x + 1)2 - 1 y = (x + 1)2 + 1

OpenStudy (anonymous):

The vertex form is this\[y = (x-h)^2 + k\]where the vertex is \((h,k)\)

OpenStudy (andertj21):

Which of the choices do i pick then?

OpenStudy (andertj21):

Well....?